OpenStudy (loser66):

Is there any short way to expand \(\dfrac{x^{1/3}-a^{1/3}}{x-a}\) Please help

OpenStudy (loser66):

@SithsAndGiggles

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \normalsize \text{let } x^{\frac{1}{3}}=m \normalsize \text{and let } a^{\frac{1}{3}}=n \hspace{.33em}\\~\\ \dfrac{x^{\frac{1}{3}}-a^{\frac{1}{3}}}{x-a} \hspace{.33em}\\~\\ =\dfrac{m-n}{m^3-n^3} \hspace{.33em}\\~\\ =\dfrac{1}{m^2+mn+n^2} \hspace{.33em}\\~\\ =\dfrac{1}{x^{\frac{2}{3}}+x^{\frac{1}{3}}a^{\frac{1}{3}}+a^{\frac{2}{3}}} \hspace{.33em}\\~\\ }\end{align}\) see if that are asking about

OpenStudy (loser66):

Thank you so much

OpenStudy (mathmath333):

was easy

OpenStudy (loser66):

Yes, for you only, not for me. I thought of polynomial expansion and got stuck. You can see what wolfram does. OMG, no way !! http://www.wolframalpha.com/input/?i=expand+%28x^%281%2F3+%29-a^%281%2F3%29%29%2F%28%28x-a%29

OpenStudy (mathmath333):

ur link is showing some series expansion

OpenStudy (loser66):

Nope, My Prof asked me to show work. hehehe...

OpenStudy (loser66):

Since it is part of the problem. I need show f(x) = x^(1/3) differentiable at x =a (a is not 0) and that f'(a) = 1/(3a^(2/3)